Technical Status 





16th to 20th August 2004

Our network will be undergoing a heavy development during the period of 16th to 20th August as part of our Science Status Capital Project.

It is anticipated that there will be no internal or external access to Email, KNect, Webmail, Intranet, Internet, Extranet, Online Handbook.

Our website will remain online throughout this period although some services may not work.

A limited service should be resumed by August 20th, with further enhancements being added throughout the remainder of August and September.

If you encounter difficulties, Please contact ICT Support via 0121 628 0010 or via e-mail.


20th April 2004 The main servers at our ISP's offices overheated and went into shutdown.  The error was caused by an engineer who incorrectly set the temperature on the air conditioning unit.  All services, including Webmail, KNect, Intranet and Website were affected.  Full service was restored within 48 hours.
3rd February 2003 An upgrade to allow users with high resolution screens to view the main menus properly caused problems with version 5.5/6 of Internet Explorer.  The fault was traced to incorrect  Javascript and corrected
8th October 2002 Our ISP has encountered problems with the Frontpage Server Extensions over the past week.  This has resulted in less updates and the inability to perform searches on the site.  Full functionality has been restored today.
18th Sept 2002 Double checking of all internal hyperlinks to maintain integrity.  3 errors found and rectified in 1940's Photo Gallery.
17th Sept 2002 Re-size of page headings and change of heading colours to blend in with page background and appear less obtrusive.
16th Sept 2002 Re-design of homepage to include less text and more photographs.